OpenStudy (anonymous):

Part 1: Explain, using complete sentences, how to simplify the expression below. Part 2: What is the simplified expression? (3x3y4)2

OpenStudy (anonymous):

is it \[(3x^3y^4)^2\]?

OpenStudy (anonymous):

YESSS

OpenStudy (anonymous):

square everything

OpenStudy (anonymous):

help pweez

OpenStudy (anonymous):

I'm confused how

OpenStudy (anonymous):

to solve this problem

OpenStudy (anonymous):

\(3^2=9, (x^3)^2=x^{2\times 3}=x^6\) and \((y^4)^2=y^{4\times 2}=y^8\)

OpenStudy (anonymous):

so when you square everything you get \[9x^6y^8\] although there is nothing inherently "simpler" about \[9x^4y^8\] than \((3x^2y^3)^2\) there is no such mathematical operation as "simplify" in this problem you are being asked to "square"

OpenStudy (anonymous):

Oh..

OpenStudy (anonymous):

thanks for the answer!